# Mind the Confidence Gap: Overconfidence, Calibration, and Distractor Effects in Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) show remarkable proficiency in natural language tasks, yet their frequent overconfidence—misalignment between predicted confidence and true correctness—poses significant risks in critical decision-making applications. We present a comprehensive analysis on calibration in LLMs across nine LLMs and three factual Question-Answering (QA) datasets, systematically comparing standard free-generation settings against structured distractor-augmented prompts. Our evaluation reveals that explicitly incorporating distractors can substantially mitigate miscalibration, achieving relative accuracy improvements up to 460% and ECE reductions up to 90%. Despite general trends, we uncover nuanced findings: large RLHF-tuned models display inherent calibration strengths but can paradoxically suffer increased miscalibration on easier queries, whereas smaller models benefit disproportionately from distractor prompts but remain significantly miscalibrated. Through detailed analyses across question types, we identify persistent calibration failures, particularly in person-based queries. We conclude with concrete recommendations—targeted fine-tuning, structured prompting, and strategic model choice—to ensure reliable, trustworthy LLM deployments. To ensure reproducibility, we provide our code and dataset [^1] https://github.com/prateekchhikara/llms-calibration.
Mind the Confidence Gap: Overconfidence, Calibration, and Distractor Effects in Large Language Models
Prateek Chhikara University of Southern California, USA pchhikar@usc.edu
@@ 1 Introduction
Large Language Models (LLMs) have significantly advanced natural language understanding, achieving state-of-the-art results across tasks including conversational AI Skjuve et al. (2024); Zhang (2024), scientific discovery Kumar (2024), and multimodal systems Zhang et al. (2023); Chhikara et al. (2024). As LLMs increasingly guide critical decisions in sensitive domains—such as healthcare, finance, and law—the reliability of their confidence estimates becomes paramount. Misalignment between model confidence and actual correctness, known as miscalibration, poses severe risks, potentially eroding user trust and causing costly or hazardous errors Dhuliawala et al. (2023); Geng et al. (2024). For example, as illustrated in Figure 1, when asked “Who received the IEEE Frank Rosenblatt Award in 2010?”, a leading LLM confidently but incorrectly answers “Geoffrey Hinton” with a confidence of 93%, despite the correct answer being “Michio Sugeno”. Such pronounced overconfidence can lead users to mistakenly trust erroneous outputs—particularly problematic in high-stakes applications such as medical diagnoses or financial decisions. Well-calibrated models, on the other hand, report confidence scores accurately reflecting their true reliability, thus enabling systems to flag uncertain predictions for human oversight and significantly mitigating real-world risks.
Modern Question-Answering pipelines adopt “one-right+several-wrong” format — whether through retrieval-augmented candidate spans, knowledge-graph sibling entities, self-consistency checks, or ensemble methods—to improve answer selection. However, these structured, distractor-rich settings introduce novel challenges for confidence estimation—challenges that classical post-hoc calibration methods (temperature scaling, Platt scaling, isotonic regression Guo et al. (2017)) were not originally designed to address. Although such methods have proven effective on small- to medium-scale neural networks, their applicability to today’s large-scale LLMs under real-world, distractor-heavy conditions remains unclear. Prior work has examined isolated factors—model scale, architecture (dense vs. mixture of experts (MoE)), and fine-tuning regime (supervised fine-tuning (SFT) vs. reinforcement learning from human feedback (RLHF)) Leng et al. (2024); Li et al. (2024)—but has not extensively examined how explicit distractors, ubiquitous in deployed QA systems, affect calibration accuracy and confidence ranking for modern LLMs.
@@
Research in cognitive psychology demonstrates that human overconfidence can be reduced by explicitly considering alternative answers before making decisions Lord et al. (1984); Mussweiler et al. (2000). Inspired by this “consider-the-opposite” strategy, we investigate whether presenting LLMs with plausible distractors similarly mitigates their systematic overconfidence and enhances calibration. To be specific, we conduct the first large-scale empirical study of LLM calibration, comparing performance under standard (free-generation) and distractor-augmented settings. Our contributions are fourfold: (1) we introduce a unified calibration benchmark evaluating nine state-of-the-art LLMs—spanning model sizes (8B–70B to greater than 1T), architectures (dense vs. MoE), and fine-tuning methods (SFT vs. RLHF)—across three factual QA datasets (SimpleQA, FaVIQ, TriviaQA); (2) we propose a structured distractor-augmented evaluation paradigm, where models select answers from one correct and multiple plausible incorrect options, enabling simultaneous assessment of accuracy improvements and Expected Calibration Error (ECE); (3) we perform fine-grained analyses across question types (e.g., person, date) to identify conditions of severe miscalibration; and (4) we systematically disentangle how scale, tuning regime, and architecture independently influence model calibration and responsiveness to distractors.
@@ 2 Related Work
Methods directly elicit uncertainty from LLMs. Prompt-based approaches that verbalize confidence Tian et al. (2023); Mielke et al. (2022); Lin et al. or aggregate multiple outputs Xiong et al. have demonstrated effectiveness, particularly for black-box models.
@@
Approaches such as calibration-aware RLHF Leng et al. (2024) and Mixup-style data augmentation Park and Caragea (2022) aim to optimize calibration during training. While these methods reduce ECE, they remain sensitive to distribution shifts Liu et al. .
@@
Chen et al. (2023) show that calibration emerges early in self-supervised pre-training, and Zhu et al. (2023) demonstrate that instruction tuning and RLHF can preserve or even enhance these gains. Jiang et al. Jiang et al. (2021) find that post-hoc methods like temperature scaling often fail to align confidence with accuracy in factual QA. However, the effects of explicit distractor-based prompting on LLM calibration remain unexplored.
@@
Techniques adjust predictions after training, with classical approaches like temperature scaling Guo et al. (2017) still underexplored for contemporary LLMs. Recent studies indicate persistent overconfidence even at larger scales, highlighting a potential degradation in calibration performance with increased model size Zhou et al. (2024a, b).
@@ 3.1 Evaluation Methods
and extract the model’s final answer $\mathbf{y}_{i}^{(D)}$ and record the associated confidence $\mathbf{c}_{i}^{(D)}$. Prompt templates $\pi_{N},\pi_{D}$ are provided in Appendix A.
3.2 Evaluation Datasets:
SimpleQA
We use the SimpleQA dataset Wei et al. (2024), which provides a reliable benchmark for evaluating LLM factual accuracy and calibration. Comprising short, fact-seeking queries with clearly defined correct answers, SimpleQA enables precise measurement of model confidence and alignment with factual correctness. Its high-quality annotations, verified by multiple independent AI trainers, ensure accuracy and unambiguity, making it well-suited for calibration assessment. The dataset contains 4326 question-answer pairs.
FaVIQ
Park et al. (2022) We select data points from test subset of the R-set. The dataset initially contains 5877 data points, out of which we focus exclusively on the 2922 data points labeled as “supports,” indicating that the provided answer is correct. FAVIQ is particularly appropriate for our experiments due to its construction methodology derived from real-world information-seeking questions. This design inherently reduces strong lexical biases found in other crowdsourced datasets, promoting nuanced semantic understanding.
TriviaQA
We use the TriviaQA dataset Joshi et al. (2017), for evaluating open-domain question answering and factual knowledge retrieval. For our experiments, we select first 1000 question-answer pairs from the validation split of the rc.web.no_content subset, ensuring a diverse yet controlled evaluation set. By restricting our selection to the no-content class, we focus on settings where models must rely purely on prior knowledge without the assistance of retrieved supporting context, isolating intrinsic model calibration behavior.
For each question-answer pair in the above mentioned datasets, we used GPT-4o-mini to generate three distractors—factually incorrect yet contextually plausible answers that matched the answer type (e.g., dates for date-based questions), remained distinct, and maintained similar specificity. Additional details on the prompt are provided in Appendix A.
@@ 3.3 Selected LLMs
We select nine representative models, spanning three major families—OpenAI’s GPT-4 series, Meta’s LLaMA lineage, and two leading open-weight models (Gemma-2 and Qwen-qwq). We select these models from OpenAI [^2] https://openai.com and GroqCloud [^3] https://groq.com API services. More details about the selected models and their taxonomy are in Appendix D.
@@
Meta’s LLaMA-3 Dubey et al. (2024) series spans three dense checkpoints: an 8B base with 8K window, an 8B “Instant” assistant-tuned model with 128K window, and a 70B base (8K window). Each uses Grouped-Query Attention (GQA) Ainslie et al. for efficient long-context processing; all these variant undergoes SFT and RLHF. In contrast, LLaMA-4-Scout-17b [^4] https://ai.meta.com/blog/llama-4-multimodal-intelligence/ adopts a 16-expert MoE transformer which supports up to 10M tokens, and is fine-tuned with both SFT and RLHF for reasoning.
@@
Gemma2-9b-it Team et al. (2024) leverages knowledge-distillation pre-training on 8T tokens—with interleaved local–global and group-query attention—followed by instruction fine-tuning (SFT + Direct Preference Optimization (DPO) Rafailov et al. (2023)) within 8K-token context. By contrast, Qwen-qwq-32b Hui et al. (2024); Yang et al. (2024) is a 32B-parameter dense model (64 layers, Rotary Positional Embedding (RoPE) Su et al. (2024), SwiGLU Dauphin et al. (2017)) pre-trained on 18T multilingual tokens without RLHF.
@@ 3.4 Evaluation Criteria
Following prior work, we use GPT-4o-mini as an LLM-based judge to classify responses as CORRECT, INCORRECT, or NOT_ATTEMPTED Packer et al. (2023); Wei et al. (2024); Chhikara et al. (2025). A response is CORRECT if it fully captures the gold target’s key information without contradiction, allowing minor variations in wording, order, or hedging. It is INCORRECT if it contains factual errors, contradictions, or misleading speculation, even if hedged. NOT_ATTEMPTED applies when a response lacks essential information without introducing errors, including vague or evasive answers. We experiment with using the same LLM for both prediction and judgment, finding that smaller LLM judges often misclassify responses or hesitate to assign NOT_ATTEMPTED when no valid answer is generated. Manual inspection confirms these issues, and further details are provided in the Appendix B.
@@ 3.5 Evaluation Metrics:
To evaluate performance, we measure correctly answered questions for both variations ($\mathcal{N}$ and $\mathcal{D}$). For calibration assessment, we use ECE to quantify the misalignment between a model’s predicted confidence and actual accuracy. A well-calibrated model produces confidence estimates that closely match its true correctness, with an ECE of zero indicating perfect calibration. Following Naeini et al. (2015), we compute ECE using empirical binning (bin size 0.1) to ensure a robust measurement of miscalibration. Additionally, we define two complementary metrics: $\mathcal{D}_{helped}$, denoting instances where the model failed under the $\mathcal{N}$ setting but succeeded when distractors were added ($\mathcal{N}_{i}=0$ and $\mathcal{D}_{i}=1$); and $\mathcal{D}_{harmed}$, capturing the reverse—cases where the model initially answered correctly under $\mathcal{N}$ but erred when distractors were introduced ($\mathcal{N}_{i}=1$ and $\mathcal{D}_{i}=0$).
@@ 4.2 Effects of Structured Distractors on Accuracy and Confidence
To quantify effect of adding correct answer alongside three incorrect options, we measure relative accuracy gain $\Delta\mathrm{Acc}=(\mathcal{D}_{\mathrm{correct}}-\mathcal{N}_{\mathrm{correct}})/\mathcal{N}_{\mathrm{correct}}$ and ECE compression $\Delta\mathrm{ECE}=\mathcal{N}_{\mathrm{ECE}}-\mathcal{D}_{\mathrm{ECE}}$. Table 1 reports counts and percentage of $\mathcal{D}_{\mathrm{helped}}$ vs. $\mathcal{D}_{\mathrm{harmed}}$ and NOT_ATTEMPTED ($\mathcal{D}_{\text{na}}$). Figure 2 overlays reliability diagrams (RDs) for all the nine models on SimpleQA dataset. RDs for FaVIQ and TriviaQA are in the Appendix E.
[…]

### v3

@@ Abstract
Large Language Models (LLMs) show remarkable proficiency in natural language tasks, yet their frequent overconfidence—misalignment between predicted confidence and true correctness—poses significant risks in critical decision-making applications. We present a comprehensive analysis on calibration in LLMs across nine LLMs and three factual Question-Answering (QA) datasets, systematically comparing standard free-generation settings against structured distractor-augmented prompts. Our evaluation reveals that explicitly incorporating distractors can substantially mitigate miscalibration, achieving relative accuracy improvements up to 460% and ECE reductions up to 90%. Despite general trends, we uncover nuanced findings: large RLHF-tuned models display inherent calibration strengths but can paradoxically suffer increased miscalibration on easier queries, whereas smaller models benefit disproportionately from distractor prompts but remain significantly miscalibrated. Through detailed analyses across question types, we identify persistent calibration failures, particularly in person-based queries. We conclude with concrete recommendations—targeted fine-tuning, structured prompting, and strategic model choice—to ensure reliable, trustworthy LLM deployments. Code is publicly available at: https://github.com/prateekchhikara/llms-calibration
@@ 1 Introduction
Large Language Models (LLMs) have significantly advanced natural language understanding, achieving state-of-the-art results across tasks including conversational AI (skjuve2024people; zhang-2024-guided), scientific discovery (kumar2024large), and multimodal systems (zhang2023visual; chhikara2024fire; zhangmllms). As LLMs increasingly guide critical decisions in sensitive domains—such as healthcare, finance, and law—the reliability of their confidence estimates becomes paramount. Misalignment between model confidence and actual correctness, known as miscalibration, poses severe risks, potentially eroding user trust and causing costly or hazardous errors (dhuliawala2023diachronic; geng2024survey). For example, as illustrated in Figure 1, when asked “Who received the IEEE Frank Rosenblatt Award in 2010?”, a leading LLM confidently but incorrectly answers “Geoffrey Hinton” with a confidence of 93%, despite the correct answer being “Michio Sugeno”. Such pronounced overconfidence can lead users to mistakenly trust erroneous outputs—particularly problematic in high-stakes applications such as medical diagnoses or financial decisions. Well-calibrated models, on the other hand, report confidence scores accurately reflecting their true reliability, thus enabling systems to flag uncertain predictions for human oversight and significantly mitigating real-world risks.
Modern Question‐Answering (QA) pipelines adopt “one‐right+several‐wrong” format — whether through retrieval‐augmented candidate spans, knowledge‐graph sibling entities, self‐consistency checks, or ensemble methods—to improve answer selection. However, these structured, distractor-rich settings introduce novel challenges for confidence estimation—challenges that classical post-hoc calibration methods (temperature scaling, Platt scaling, isotonic regression (guo2017calibration)) were not originally designed to address. Although such methods have proven effective on small‐ to medium‐scale neural networks, their applicability to today’s large‐scale LLMs under real-world, distractor-heavy conditions remains unclear. Prior work has examined isolated factors—model scale, architecture (dense vs. mixture of experts (MoE)), and fine-tuning regime (supervised fine-tuning (SFT) vs. reinforcement learning from human feedback (RLHF)) (leng2024taming; li2024can)—but has not extensively examined how explicit distractors, ubiquitous in deployed QA systems, affect calibration accuracy and confidence ranking for modern LLMs.
@@
Research in cognitive psychology demonstrates that human overconfidence can be reduced by explicitly considering alternative answers before making decisions (lord1984considering; mussweiler2000overcoming). Inspired by this “consider-the-opposite” strategy, we investigate whether presenting LLMs with plausible distractors similarly mitigates their systematic overconfidence and enhances calibration. To be specific, we conduct the first large-scale empirical study of LLM calibration, comparing performance under standard (free-generation) and distractor-augmented settings. Our contributions are fourfold: (1) we introduce a unified calibration benchmark evaluating nine state-of-the-art LLMs—spanning model sizes (8B–70B to greater than 1T), architectures (dense vs. MoE), and fine-tuning methods (SFT vs. RLHF)—across three factual QA datasets (SimpleQA, FaVIQ, TriviaQA); (2) we propose a structured distractor-augmented evaluation paradigm, where models select answers from one correct and multiple plausible incorrect options, enabling simultaneous assessment of accuracy improvements and Expected Calibration Error (ECE); (3) we perform fine-grained analyses across question types (e.g., person, date) to identify conditions of severe miscalibration; and (4) we systematically disentangle how scale, tuning regime, and architecture independently influence model calibration and responsiveness to distractors.
@@ 2 Related Work
Methods directly elicit uncertainty from LLMs. Prompt-based approaches that verbalize confidence (tian2023just; mielke2022reducing; linteaching) or aggregate multiple outputs (xiongcan) have demonstrated effectiveness, particularly for black-box models.
@@
Approaches such as calibration-aware RLHF (leng2024taming) and Mixup-style data augmentation (park2022calibration) aim to optimize calibration during training. While these methods reduce ECE, they remain sensitive to distribution shifts (liucalibration).
@@
chen2023close show that calibration emerges early in self-supervised pre-training, and zhu-etal-2023-calibration demonstrate that instruction tuning and RLHF can preserve or even enhance these gains. jiang2021can find that post-hoc methods like temperature scaling often fail to align confidence with accuracy in factual QA. However, the effects of explicit distractor-based prompting on LLM calibration remain unexplored.
@@
Techniques adjust predictions after training, with classical approaches like temperature scaling (guo2017calibration) still underexplored for contemporary LLMs. Recent studies indicate persistent overconfidence even at larger scales, highlighting a potential degradation in calibration performance with increased model size (zhou2024relying; zhou2024larger).
@@ 3 Experimental Setup
3.1 Evaluation Datasets:
SimpleQA
We use the SimpleQA dataset (wei2024measuring), which provides a reliable benchmark for evaluating LLM factual accuracy and calibration. Comprising short, fact-seeking queries with clearly defined correct answers, SimpleQA enables precise measurement of model confidence and alignment with factual correctness. Its high-quality annotations, verified by multiple independent AI trainers, ensure accuracy and unambiguity, making it well-suited for calibration assessment. The dataset contains 4326 question-answer pairs.
FaVIQ
(park2022faviq) We select data points from test subset of the R-set. The dataset initially contains 5877 data points, out of which we focus exclusively on the 2922 data points labeled as “supports,” indicating that the provided answer is correct. FaVIQ is particularly appropriate for our experiments due to its construction methodology derived from real-world information-seeking questions. This design inherently reduces strong lexical biases found in other crowdsourced datasets, promoting nuanced semantic understanding.
TriviaQA
We use the TriviaQA dataset (joshi2017triviaqa), for evaluating open-domain question answering and factual knowledge retrieval. For our experiments, we select first 1000 question-answer pairs from the validation split of the rc.web.no_content subset, ensuring a diverse yet controlled evaluation set. By restricting our selection to the no-content class, we focus on settings where models must rely purely on prior knowledge without the assistance of retrieved supporting context, isolating intrinsic model calibration behavior.
@@ 3.2 Evaluation Methods
and extract the model’s final answer $\mathbf{y}_{i}^{(D)}$ and record the associated confidence $\mathbf{c}_{i}^{(D)}$. Prompt templates $\pi_{N},\pi_{D}$ are provided in Appendix A. Our distractors were generated using GPT-4o-mini with a carefully designed prompt to ensure they were factually incorrect yet contextually plausible. Specifically, for each question–answer pair in the datasets, we used GPT-4o-mini to generate three distractors that (i) matched the expected answer type (e.g., dates for date-based questions), (ii) remained distinct from the correct answer, and (iii) maintained comparable specificity and context. To further validate plausibility, we manually inspected over 500 randomly sampled examples from the SimpleQA dataset, confirming that the generated distractors were consistently plausible and contextually relevant.
Why Elicited Confidence?
We measure confidence via elicited self-reports (0–100) because they capture task-level belief (“How sure are you that your answer is correct?”) rather than token-level fluency artifacts. Prior work shows that verbalized probabilities can better reflect correctness than raw token likelihoods, which are sensitive to phrasing and tokenization (linteaching). In RLHF-tuned systems, elicited confidence has also been observed to track calibration more reliably than log-probabilities, which can degrade post-alignment (tian2023just). This aligns with “linguistic calibration” evidence that making models state their confidence reduces overconfidence and improves user-facing transparency (mielke2022reducing). Elicitation is also the only uniformly available signal across the nine black-box and open-weight models we evaluate, enabling apples-to-apples comparison. We do not claim it is the sole or optimal measure: logit-based margins/entropy and self-consistency (majority-vote variance) are complementary signals. Our findings should therefore be interpreted in the context of elicited confidence; we add this clarification to promote comparability and to reflect the literature’s guidance on when verbalized uncertainty is informative.
@@ 3.3 Selected LLMs
We select nine representative models, spanning three major families—OpenAI’s GPT-4 series, Meta’s LLaMA lineage, and two leading open‑weight models (Gemma-2 and Qwen-qwq). We select these models from OpenAI [^1] https://openai.com and GroqCloud [^2] https://groq.com API services. More details about the selected models and their taxonomy are in Table 1.
GPT-4o
(undisclosed / 128K)
GPT-4o-mini
(undisclosed / 128K)
Dense Multi-modal (web text, code, images) SFT + RLHF
GPT-4-turbo
(undisclosed / 128K)
Dense Multi-modal (web text, code, images) Pre-training + SFT + RLHF
LLaMA-3.1-8B-Instant
(8B / 128K)
LLaMA-3-8B-Instruct
(8B / 8K)
LLaMA-3-70B-Instruct
(70B / 8K)
LLaMA-4-Scout-17B
(17B / 10M)
Gemma2-9B-it
(9B / 8K)
Qwen-qwq-32B
(32B / 131K)
@@
Meta’s LLaMA-3 (dubey2024llama) series spans three dense checkpoints: an 8B base with 8K window, an 8B “Instant” assistant-tuned model with 128K window, and a 70B base (8K window). Each uses Grouped-Query Attention (GQA) (ainslie2023gqa) for efficient long-context processing; all these variant undergoes SFT and RLHF. In contrast, LLaMA-4-Scout-17b [^3] https://ai.meta.com/blog/llama-4-multimodal-intelligence/ adopts a 16-expert MoE transformer which supports up to 10M tokens, and is fine-tuned with both SFT and RLHF for reasoning.
@@
Gemma2-9b-it (team2024gemma) leverages knowledge-distillation pre-training on 8T tokens—with interleaved local–global and group-query attention—followed by instruction fine-tuning (SFT + Direct Preference Optimization (DPO) (rafailov2023direct)) within 8K-token context. By contrast, Qwen-qwq-32b (hui2024qwen2; yang2024qwen2) is a 32B-parameter dense model (64 layers, Rotary Positional Embedding (RoPE) (su2024roformer), SwiGLU (dauphin2017language)) pre-trained on 18T multilingual tokens without RLHF.
The GPT-4 family and LLaMA-3 series are dense models trained on massive multimodal or text‐only corpora, each undergoing both SFT and RLHF before instruct-tuning. In contrast, LLaMA-4-Scout employs a 16-expert MoE design over mixed vision-text data, while Gemma2-9b-it and Qwen-qwq-32b explore distillation-based and pure SFT regimes, respectively. Despite these varied strategies, all nine models receive dedicated fine-tuning and instruction-tuning to optimize performance and calibration in downstream QA and conversational tasks.
@@ 3.4 Evaluation Criteria
Following prior work, we use GPT-4o-mini as an LLM-based judge to classify responses as CORRECT, INCORRECT, or NOT_ATTEMPTED (packer2023memgpt; wei2024measuring; chhikara2025mem0). A response is CORRECT if it fully captures the gold target’s key information without contradiction, allowing minor variations in wording, order, or hedging. It is INCORRECT if it contains factual errors, contradictions, or misleading speculation, even if hedged. NOT_ATTEMPTED applies when a response lacks essential information without introducing errors, including vague or evasive answers. We experiment with using the same LLM for both prediction and judgment, finding that smaller LLM judges often misclassify responses or hesitate to assign NOT_ATTEMPTED when no valid answer is generated. Manual inspection confirms these issues, and further details are provided in the Appendix B.
@@ 3.5 Evaluation Metrics:
To evaluate performance, we measure correctly answered questions for both variations ($\mathcal{N}$ and $\mathcal{D}$). For calibration assessment, we use ECE to quantify the misalignment between a model’s predicted confidence and actual accuracy. A well-calibrated model produces confidence estimates that closely match its true correctness, with an ECE of zero indicating perfect calibration. Following (naeini2015obtaining), we compute ECE using empirical binning (bin size 0.1) to ensure a robust measurement of miscalibration. Additionally, we define two complementary metrics: $\mathcal{D}_{helped}$, denoting instances where the model failed under the $\mathcal{N}$ setting but succeeded when distractors were added ($\mathcal{N}_{i}=0$ and $\mathcal{D}_{i}=1$); and $\mathcal{D}_{harmed}$, capturing the reverse—cases where the model initially answered correctly under $\mathcal{N}$ but erred when distractors were introduced ($\mathcal{N}_{i}=1$ and $\mathcal{D}_{i}=0$).
@@ 4.2 Effects of Structured Distractors on Accuracy and Confidence
To quantify effect of adding correct answer alongside three incorrect options, we measure relative accuracy gain $\Delta\mathrm{Acc}=(\mathcal{D}_{\mathrm{correct}}-\mathcal{N}_{\mathrm{correct}})/\mathcal{N}_{\mathrm{correct}}$ and ECE compression $\Delta\mathrm{ECE}=\mathcal{N}_{\mathrm{ECE}}-\mathcal{D}_{\mathrm{ECE}}$. Table 2 reports counts and percentage of $\mathcal{D}_{\mathrm{helped}}$ vs. $\mathcal{D}_{\mathrm{harmed}}$ and NOT_ATTEMPTED ($\mathcal{D}_{\text{na}}$). Figure 2 overlays reliability diagrams (RDs) for all the nine models on SimpleQA dataset. RDs for FaVIQ and TriviaQA are in the Appendix C.
Across all models, structured distractors boost accuracy and (for most cases) reduce ECE. On the challenging SimpleQA benchmark, distractors often halve ECE (up to $\Delta\mathrm{ECE}\approx 0.4$) and more than double accuracy for smaller variants, confirming that explicit options help recalibrate confidence when generation alone is unreliable. However, on TriviaQA—the easiest dataset—GPT-4o and GPT-4-turbo (the largest models) exhibit a slight increase in ECE under $\mathcal{D}$, despite relative accuracy gains of 3–5%. We observe that $\mathrm{ECE}_{\mathcal{N}\to\mathcal{D}}(\texttt{GPT-4o})$ rises from 0.071 to 0.083, and $\mathrm{ECE}_{\mathcal{N}\to\mathcal{D}}(\texttt{GPT-4-turbo})$ rises from 0.025 to 0.048. This counterintuitive effect likely stems from confidence inflation: on already-easy examples, the multiple-choice context encourages the model to assign excessive probability mass to the correct answer, amplifying residual misalignment between predicted confidence and true correctness.